### v1

# Exponentially Enhanced Scheme for the Heralded Qudit GHZ State in Linear Optics

Abstract
High-dimensional multipartite entanglement plays a crucial role in quantum information science. However, existing schemes for generating such entanglement become increasingly complex and costly as the dimension of quantum units increases. In this work, we overcome the limitation by proposing a significantly enhanced linear optical heralded scheme that generates the $d$-level $N$-partite GHZ state with single-photon sources and their linear operations. Our scheme requires $dN$ photons to generate the target state with substantially improved success probability from previous schemes. It employs linear optical logic gates compatible with any qudit encoding system and can generate generalized GHZ states with installments of beamsplitters. With efficient generations of high-dimensional resource states, our work opens avenues for further exploration in high-dimensional quantum information processing.
Introduction.—
A qubit, which is comprised of a two-level quantum system and functions, is the quantum counterpart to the classical bit. It is generally considered the basic unit that stores and delivers information in quantum tasks. However, many natural systems have higher degrees of freedom, such as frequencies and paths of particles. Quantum information theory provides a new way of encoding such multi-level systems, which is called the qudit. Qudits provide larger Hilbert spaces than qubits for encoding and manipulating information. Higher-dimensional quantum systems exhibit distinctive properties that are unattainable with qubit-based systems such as contextuality [1]. They also provide a handy platform to simulate several quantum systems [2, 3].
Numerous advantages arise from entangling multiple quantum units in quantum information processing. The Bell state of qubits facilitates quantum enhancements over classical counterparts [4, 5]. Beyond two-qubit systems, quantum entanglement in multipartite and/or higher-dimensional systems is essential for new quantum technologies such as quantum cryptography, reduction of communication complexity, quantum communication protocols, and random number generation [6, 7, 8, 9, 10].
On the other hand, the intricate structure of qudit multipartite entanglement brings about several theoretical and practical challenges such as classification problems according the separability of states [11] and the maximal quantum violation by non-maximally entangled qutrit bipartite states [12, 13]. Another crucial problem lies on the generation of qudit mulipartite entangled states for the practical performances of quantum tasks. The increased degrees of freedom in the system generally make the generation process significantly more complex and expensive.
We contribute to overcoming this limit by providing a magnificently enhanced heralded scheme for generating the qudit multipartite entanglement in linear optical networks (LONs). The LON is one of promising systems that can feasibly generate quantum resource states. Non-destructive entanglement generations by heralding are possible in LONs by utilizing ancillary particles to herald the successful operation events without direct detections of the target states. Therefore, photons entangled by heralding become proper resources for quantum computations with multiple gates [14, 15] and also facilitates loophole-free Bell tests [16, 17]. However, the number of additional particles and modes that the heralded schemes require increases as the dimension of entanglement becomes higher, which makes the circuit structures much more intricate. Therefore, it is prerequisite to design optimized circuits that achieve the target state with minimal resources for the efficient high-dimensional quantum information processing.
In this work, we propose a heralded scheme to generate one of essential qudit mutipartite entangled states, i.e., the $N$-partite $d$-level GHZ state $|GHZ_{N}^{d}\rangle$:
$|GHZ_{N}^{d}\rangle=\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}|j\rangle_{1}\otimes\cdots\otimes|j\rangle_{N}$ (1)
where $|j\rangle_{k}$ denotes the $k$th qudit being in the $j$th state. The qudit GHZ state reveals intriguing non-classical properties of $N$-partite $d$-level systems such as the GHZ paradox [18, 19, 20, 21] and Bell inequalities [22]. It also becomes a resource state to generate qudit cluster states for the high-dimensional measurement-based quantum computing (MBQC) [23, 24, 25].
Heralded schemes for the qudit GHZ state in various physical setups have been proposed, e.g., with single-photon sources and linear optical operators [25], EPR sources [26], or an array of non-interacting photon emitters [27]. Among them, Ref. [25] suggested a linear optical heralded scheme based on the zero-transmission laws (ZTL) in quantum Fourier tranforms (QFTs). However, it suffers from low efficiency and exponentially increasing photon numbers as $N$ and $d$ grows.
In this work, we propose a scheme that generates the GHZ state of any $N$ and $d$ with the most feasible optical elements, i.e., single photons and linear optical operators as in Ref. [25]. Our scheme has several manifest advantages over formerly suggested schemes: First, it is an efficient circuit with significantly less photons ($dN$ photons for $|GHZ_{N}^{d}\rangle$) and exponentially higher success probability than that in Ref. [25]. Second, it can easily control the probability amplitudes of each term in the GHZ state by simply placing beamsplitters on the paths of photons, which result in the generation of the generalized qudit GHZ state. Third, it consists of linear optical logic gates that any type of qubit encoding (e.g., optical angular momentum (OAM), multi-rail, time-bin, etc.) can establish.
Linear optical logic gates for generating $|GHZ_{N}^{d}\rangle$.—
We first introduce a set of linear optical logic gates, which will compose our scheme. The list of the gates with their pictorial notations is given in Fig. 1.
Figure 1: The list of linear optical logic gates that will be used to design the qudit GHZ generating circuits.
1. $CN_{d}$: The $d$-level generalized CNOT gate between the spatial mode $j$ ($(\in\{0,1,\cdots,d\})$ and the internal state $\tilde{s}$ ($(\in\{\tilde{0},\tilde{1},\cdots,\tilde{d}\})$, which works on states as
$|\tilde{s}\rangle_{j}={\hat{a}}^{\dagger}_{j,\tilde{s}}|vac\rangle$
$\to$ $~{}CN_{d}|\tilde{s}\rangle_{j}=|\tilde{s}\rangle_{j\oplus_{d}\tilde{s}}={\hat{a}}^{\dagger}_{j\oplus_{d}\tilde{s},\tilde{s}}|vac\rangle$ (2)
(${\hat{a}^{\dagger}}_{j,\tilde{s}}$ is a photon creation operator, $|vac\rangle$ is the vacuum state, and $\oplus_{d}$ is the addition modular $d$). Its inverse $\overline{CN_{d}}$ works as
$|\tilde{s}\rangle_{j}={\hat{a}}^{\dagger}_{j,\tilde{s}}|vac\rangle$
$\to$ $~{}\overline{CN}_{d}|\tilde{s}\rangle_{j}=|\tilde{s}\rangle_{j\ominus_{d}\tilde{s}}={\hat{a}}^{\dagger}_{j\ominus_{d}\tilde{s},\tilde{s}}|vac\rangle$ (3)
where dsubscriptsymmetric-$\ominus_{d}$ is the subtraction modular $d$. These operators split states in the same spatial mode with different internal states into different spatial modes (see Appendix A for a more detailed explanation).
For $d=2$, we can build these operators with polarized photons. By setting $\{H,V\}=\{0,1\}$ ($H$=horizontal, $V$=vertical), $CN_{2}$ and $\overline{CN}_{2}$ are both implemented with a polarizing beam splitter (PBS).
2. $F_{d}$: The internal state Fourier-transformation gate,
$F_{d}|\tilde{s}\rangle_{j}=|s\rangle_{j}=\frac{1}{\sqrt{d}}\sum_{r=0}^{d-1}e^{i\omega r}|\tilde{r}\rangle~{}~{}~{}(\omega=e^{i\frac{2\pi k}{d}}).$ (4)
$F_{d}$ changes the computational basis of the internal state into its Fourier-transformed basis. For $d=2$, $F_{2}$ can be installed in the polarization encoding with a half-wave plate (HWP).
3. $X_{d}$: The $d$-level unitary generalization of $\sigma_{x}$ in qubit,
$X_{d}=\sum_{k=0}^{d-1}|k\oplus 1\rangle\langle k|,$ (5)
under which $|k\rangle$ transforms to $|k\oplus 1\rangle$.
4. $BS_{t}$: A beamsplitter that transmits a photon with transitivity $t$.
5. $D_{F}$: Number resolving detections in the Fourier-transformed basis, i.e., measurements in $\{|s\rangle\}_{s=0}^{d-1}$ and postselections of the cases when only one particle arrives at the detector. This is achieved by combining $F_{d}$ and $CN_{d}$, i.e.,
where [Uncaptioned image][Uncaptioned $\end{gathered}$ denotes the particle-number-resolving detector.
Linear optical circuits for generating $|GHZ_{N}^{d}\rangle$.—
By combining the gates listed in Fig. 1, we can design a circuit that generates $|GHZ_{N}^{d}\rangle$ for any $N$ and $d$. Our scheme is given in Fig. 2, which can be considered a linear optical installation of the boson subtraction scheme in Ref. [28] to generate the same target state. While Ref. [28] lacks how to furnish the qudit boson subtraction operators with real physical systems, we demonstrate that the elements of the boson subtraction scheme for $|GHZ_{N}^{d}\rangle$ can be translated to the linear optical logic gates to construct a physical qudit GHZ scheme with heralding. A more detailed explanation on the relation between the subtraction operator and our optical heralded scheme is given in Appendix B.
In each input mode of the circuit in Fig. 2, we inject $d$ photons with different internal states in the same mode. The operation of each dashed box in the circuit corresponds to the boson subtraction of $(d-1)$ particles, hence $(d-1)N$ subtraction of photons in total. $BS_{t}$s are attached to control the probability amplitude of the target state, hence we can obtain different states in the class of generalized GHZ states by changing the transitivity of the $BS_{t}$s. It is worth noting that our circuit inherits the cyclic symmetry of the qudit GHZ state.
The success probability $P_{suc}(N,d)$ for generating $|GHZ_{N}^{d}\rangle$ is given by
$P_{suc}(N,d)=d\Big{(}\frac{d!}{d^{d}}\Big{)}^{2N}.$ (6)
When $d=2$, by considering $CN_{2}$ and $\overline{CN}_{2}$ with PBS and $F_{d}$ with HWP, this circuit corresponds to the qubit GHZ generation scheme proposed in Ref. [29]. Appendix B demonstrates how the circuit in Fig. 2 generates $|GHZ_{N}^{d}\rangle$ by heralding.
Figure 2: The linear optical circuit that generates $|GHZ_{N}^{d}\rangle.$ ABSs are attached at the end of the $d$ sub-modes of the first spatial subsystem $1$ to control the probability amplitudes of the terms in the target state, hence we can obtain the generalized qudit GHZ states.
$|GHZ_{3}^{3}\rangle$ example.—
As proof of concept, we provide here a step-by-step analysis on the $N=d=3$ example (Fig. 3). There are three spatial subsystems $\{a,b,c\}$ that has three sub-modes respectively, hence the photon creation operator is denoted as $\{{\hat{a}^{\dagger}}_{j,s},{\hat{b}^{\dagger}}_{k,r},{\hat{c}^{\dagger}}_{l,t}\}$ where $\{j,k,l\}$ denotes the spatial sub-modes and $\{s,r,t\}$ the qudit states ($j,k,l,s,r,t\in\{0,1,2\}$).
Figure 3: $|GHZ_{3}^{3}\rangle$ example of our scheme. We need a BS with transitivity $\frac{1}{\sqrt{3}}$ on $X$, $Y$ and $Z$ to carve the extra amplitude of $|111\rangle$.
Step 1. Initial state preparation: We inject three photons of orthogonal internal states with each other into the 0th modes of $a$, $b$, and $c$ respectively, hence the initial state is prepared as
${\hat{a}}^{\dagger}_{0,0}{\hat{a}}^{\dagger}_{0,1}{\hat{a}}^{\dagger}_{0,2}{\hat{b}}^{\dagger}_{0,0}{\hat{b}}^{\dagger}_{0,1}{\hat{b}}^{\dagger}_{0,2}\hat{c}^{\dagger}_{0,0}\hat{c}^{\dagger}_{0,1}\hat{c}^{\dagger}_{0,2}|vac\rangle.$ (7)
Step 2. $CN_{3}$ operation: This operator works on the basis of $\{|\tilde{j}\rangle\}_{j=0}^{2}$, hence the initial state is transformed as
$\frac{1}{\sqrt{3}^{9}}$ $\big{(}{\hat{a}}^{\dagger 3}_{0,\tilde{0}}+{\hat{a}}^{\dagger 3}_{1,\tilde{1}}+{\hat{a}}^{\dagger 3}_{2,\tilde{2}}-3{\hat{a}}^{\dagger}_{0,\tilde{0}}{\hat{a}}^{\dagger}_{1,\tilde{1}}{\hat{a}}^{\dagger}_{2,\tilde{2}}\big{)}$
×$\times\big{(}{\hat{b}}^{\dagger 3}_{0,\tilde{0}}+{\hat{b}}^{\dagger 3}_{1,\tilde{1}}+{\hat{b}}^{\dagger 3}_{2,\tilde{2}}-3{\hat{b}}^{\dagger}_{0,\tilde{0}}{\hat{b}}^{\dagger}_{1,\tilde{1}}{\hat{b}}^{\dagger}_{2,\tilde{2}}\big{)}$
×$\times\big{(}\hat{c}^{\dagger 3}_{0,\tilde{0}}+\hat{c}^{\dagger 3}_{1,\tilde{1}}+\hat{c}^{\dagger 3}_{2,\tilde{2}}-3\hat{c}^{\dagger}_{0,\tilde{0}}\hat{c}^{\dagger}_{1,\tilde{1}}\hat{c}^{\dagger}_{2,\tilde{2}}\big{)}|vac\rangle$ (8)
Step 3. Rewiring: Now we rewire so that creation operators $\{{\hat{a}^{\dagger}}_{1,\tilde{1}},{\hat{b}^{\dagger}}_{1,\tilde{1}},{\hat{c}^{\dagger}}_{1,\tilde{1}}\}$ are sent to $\{\hat{X}^{\dagger}_{\tilde{1}},\hat{Y}^{\dagger}_{\tilde{1}},\hat{Z}^{\dagger}_{\tilde{1}}\}$ and $\{{\hat{a}^{\dagger}}_{2,\tilde{2}},{\hat{b}^{\dagger}}_{2,\tilde{2}},{\hat{c}^{\dagger}}_{2,\tilde{2}}\}$ to $\{{\hat{b}^{\dagger}}_{2,\tilde{2}},{\hat{c}^{\dagger}}_{2,\tilde{2}},{\hat{a}^{\dagger}}_{2,\tilde{2}}\}$ :
$\frac{1}{\sqrt{3}^{9}}$ $\big{(}{\hat{a}}^{\dagger 3}_{0,\tilde{0}}+\hat{X}^{\dagger 3}_{\tilde{1}}+{\hat{b}}^{\dagger 3}_{2,\tilde{2}}-3{\hat{a}}^{\dagger}_{0,\tilde{0}}\hat{X}^{\dagger}_{\tilde{1}}{\hat{b}}^{\dagger}_{2,\tilde{2}}\big{)}$
×$\times\big{(}{\hat{b}}^{\dagger 3}_{0,\tilde{0}}+\hat{Y}^{\dagger 3}_{\tilde{1}}+\hat{c}^{\dagger 3}_{2,\tilde{2}}-3{\hat{b}}^{\dagger}_{0,\tilde{0}}\hat{Z}^{\dagger}_{\tilde{1}}\hat{c}^{\dagger}_{2,\tilde{2}}\big{)}$
×$\times\big{(}\hat{c}^{\dagger 3}_{0,\tilde{0}}+\hat{Z}^{\dagger 3}_{\tilde{1}}+\hat{a}^{\dagger 3}_{2,\tilde{2}}-3\hat{c}^{\dagger}_{0,\tilde{0}}\hat{X}^{\dagger}_{\tilde{1}}\hat{a}^{\dagger}_{2,\tilde{2}}\big{)}|vac\rangle$ (9)
Step 4. $X_{d}$ and $F_{3}$: Now the internal states of the creation operators go back to the basis of $\{|j\rangle\}_{j=0}^{2}$.
$\frac{1}{\sqrt{3}^{9}}$ $\big{(}{\hat{a}}^{\dagger 3}_{0,0}+\hat{X}^{\dagger 3}_{\tilde{1}}+{\hat{b}}^{\dagger 3}_{2,0}-3{\hat{a}}^{\dagger}_{0,0}\hat{X}^{\dagger}_{\tilde{1}}{\hat{b}}^{\dagger}_{2,0}\big{)}$
$\times$ $\big{(}{\hat{b}}^{\dagger 3}_{0,0}+\hat{Y}^{\dagger 3}_{\tilde{1}}+{\hat{c}}^{\dagger 3}_{2,0}-3{\hat{b}}^{\dagger}_{0,0}\hat{Y}^{\dagger}_{\tilde{1}}{\hat{c}}^{\dagger}_{2,0}\big{)}$
$\times$ $\big{(}\hat{c}^{\dagger 3}_{0,0}+\hat{Z}^{\dagger 3}_{\tilde{1}}+{\hat{a}}^{\dagger 3}_{2,0}-3\hat{c}^{\dagger}_{0,0}\hat{Z}^{\dagger}_{\tilde{1}}{\hat{a}}^{\dagger}_{2,0}\big{)}|vac\rangle$ (10)
Step 5. $\overline{CN_{3}}$
$\frac{1}{3^{9}}\Big{[}({\hat{a}^{\dagger}}_{0,\tilde{0}}+{\hat{a}^{\dagger}}_{2,\tilde{1}}+{\hat{a}^{\dagger}}_{1,\tilde{2}})^{3}+\hat{X}^{\dagger 3}_{\tilde{1}}+({\hat{b}^{\dagger}}_{2,\tilde{0}}+{\hat{b}^{\dagger}}_{1,\tilde{1}}+{\hat{b}^{\dagger}}_{0,\tilde{2}})^{3}$
$~{}~{}~{}-3\sqrt{3}\alpha({\hat{a}^{\dagger}}_{0,\tilde{0}}+{\hat{a}^{\dagger}}_{2,\tilde{1}}+{\hat{a}^{\dagger}}_{1,\tilde{2}})\hat{X}^{\dagger}_{\tilde{1}}({\hat{b}^{\dagger}}_{2,\tilde{0}}+{\hat{b}^{\dagger}}_{1,\tilde{1}}+{\hat{b}^{\dagger}}_{0,\tilde{2}})\Big{]}$
×$~{}\times\Big{[}({\hat{b}^{\dagger}}_{0,\tilde{0}}+{\hat{b}^{\dagger}}_{2,\tilde{1}}+{\hat{b}^{\dagger}}_{1,\tilde{2}})^{3}+\hat{Y}^{\dagger 3}_{\tilde{1}}+({\hat{c}^{\dagger}}_{2,\tilde{0}}+{\hat{c}^{\dagger}}_{1,\tilde{1}}+{\hat{c}^{\dagger}}_{0,\tilde{2}})^{3}$
$~{}~{}~{}-3\sqrt{3}\alpha({\hat{b}^{\dagger}}_{0,\tilde{0}}+{\hat{b}^{\dagger}}_{2,\tilde{1}}+{\hat{b}^{\dagger}}_{1,\tilde{2}})\hat{Y}^{\dagger}_{\tilde{1}}({\hat{c}^{\dagger}}_{2,\tilde{0}}+{\hat{c}^{\dagger}}_{1,\tilde{1}}+{\hat{c}^{\dagger}}_{0,\tilde{2}})\Big{]}$
×$~{}\times\Big{[}({\hat{c}^{\dagger}}_{0,\tilde{0}}+{\hat{c}^{\dagger}}_{2,\tilde{1}}+{\hat{c}^{\dagger}}_{1,\tilde{2}})^{3}+\hat{Z}^{\dagger 3}_{\tilde{1}}+({\hat{a}^{\dagger}}_{2,\tilde{0}}+{\hat{a}^{\dagger}}_{1,\tilde{1}}+{\hat{a}^{\dagger}}_{0,\tilde{2}})^{3}$
$~{}~{}~{}-3\sqrt{3}\alpha({\hat{c}^{\dagger}}_{0,\tilde{0}}+{\hat{c}^{\dagger}}_{2,\tilde{1}}+{\hat{c}^{\dagger}}_{1,\tilde{2}})\hat{Z}^{\dagger}_{\tilde{1}}({\hat{a}^{\dagger}}_{2,\tilde{0}}+{\hat{a}^{\dagger}}_{1,\tilde{1}}+{\hat{a}^{\dagger}}_{0,\tilde{2}})\Big{]}|vac\rangle$ (11)
Step 6. Postselections and feed-forward: After the postselection of the cases when each detector observes only one photon, the terms that contribute to the final states are
$\frac{1}{3^{9}}\Big{[}$ $6^{3}{\hat{a}^{\dagger}}_{0\tilde{0}}{\hat{a}^{\dagger}}_{2\tilde{1}}{\hat{a}^{\dagger}}_{1\tilde{2}}{\hat{b}^{\dagger}}_{0\tilde{0}}{\hat{b}^{\dagger}}_{2\tilde{1}}{\hat{b}^{\dagger}}_{1\tilde{2}}{\hat{c}^{\dagger}}_{0\tilde{0}}{\hat{c}^{\dagger}}_{2\tilde{1}}{\hat{c}^{\dagger}}_{1\tilde{2}}$
$+6^{3}{\hat{a}^{\dagger}}_{2\tilde{0}}{\hat{a}^{\dagger}}_{1\tilde{1}}{\hat{a}^{\dagger}}_{0\tilde{2}}{\hat{b}}^{\dagger}_{2\tilde{0}}{\hat{b}}^{\dagger}_{1\tilde{1}}{\hat{b}}^{\dagger}_{0\tilde{2}}{\hat{c}^{\dagger}}_{2\tilde{0}}{\hat{c}^{\dagger}}_{1\tilde{1}}{\hat{c}^{\dagger}}_{0\tilde{2}}$
$-(3\sqrt{3})^{3}\alpha^{3}\hat{X}_{\tilde{1}}\hat{Y}_{\tilde{1}}\hat{Z}_{\tilde{1}}({\hat{a}^{\dagger}}_{1,\tilde{1}}{\hat{a}^{\dagger}}_{2,\tilde{1}}+{\hat{a}^{\dagger}}_{1,\tilde{2}}{\hat{a}^{\dagger}}_{2,\tilde{0}})$
×$~{}~{}~{}\times({\hat{b}^{\dagger}}_{1,\tilde{1}}{\hat{b}^{\dagger}}_{2,\tilde{1}}+{\hat{b}^{\dagger}}_{1,\tilde{2}}{\hat{b}^{\dagger}}_{2,\tilde{0}})({\hat{c}^{\dagger}}_{1,\tilde{1}}{\hat{c}^{\dagger}}_{2,\tilde{1}}+{\hat{c}^{\dagger}}_{1,\tilde{2}}{\hat{c}^{\dagger}}_{2,\tilde{0}})\Big{]}|vac\rangle$ (12)
After the postselection with $D_{F}$s and feedforwards, the final state becomes
$\frac{6^{3}}{3^{9}}\Big{[}|0,0,0\rangle-3\sqrt{3}\alpha^{3}|1,1,1\rangle+|2,2,2\rangle\Big{]}.$ (13)
Setting $\alpha=\frac{1}{\sqrt{3}}$, the final state becomes the tripartite qutrit GHZ state. The success probability is given by
$P_{suc}(3,3)=(\frac{2}{9})^{6}\sim 3.6\times 10^{-4}.$ (14)
Comparison with the linear optical schemes in Ref. [25].—
Compared to the schemes using the zero-transmission law (ZTL) in Fourier transform interferometers in Ref. [25] (which we call ‘ZTL scheme’ from now on), our scheme is substantially more efficient for $N\geq 3$ both in the photon numbers and success probabilities. For $N=2$ (qudit Bell state), our scheme has lower success probability ($d\Big{(}\frac{d!}{d^{d}}\Big{)}^{4}$ is smaller than $\frac{d(2d-1)!}{(2d+1)^{2d-1}}$ of the ZTL scheme) but requires less photons (we need $2d$ photons while the ZTL scheme needs $12d+$ 1).
Since the general ZTL scheme in Ref. [25] has a very low efficiency, an alternative near optimal scheme is proposed that can be applied to general $N$ and $d$. We can see that the near-optimal solution still needs significantly more photons and exponentially lower success probability than ours. Fig. 4 compares the required photon numbers and $P_{suc}$ of the near-optimal ZTL scheme with ours for $N=3$ and $2\leq d\leq 8$. On the other hand, a brute-fore method is also used in Ref. [25] for $N=d=3$ that diminishes the photon number to $25$ with $P_{suc}\sim 10^{-10}$, which is still much less efficient than our scheme that requires $9$ photons with $P_{suc}\sim 3.6\times 10^{-4}$. The authors increased $P_{suc}$ to $\sim 0.8\times 10^{-4}$ with multiplexing, which however is still lower than a quarter of ours.
Figure 4: Comparison of our scheme with the near-optimal scheme of Ref. [25] for $N=3$ in (a) photon numbers, and (b) the logarithmic scales of the success probabilities. Graph (b) shows that the success probability is enhance exponentially in our scheme.
Construction of linear optical schemes in OAM and multi-rail encodings.—
The linear optical logic gates in Fig. 1 can be implemented with any qudit encoding system of linear optics. Among them, we discuss the OAM and multi-rail encodings here.
In the OAM encoding, it is straightforward to establish the scheme, because the the OAM beam splitter [30] and OAM-only Fourier transformation operators [31] implement the $CN_{d}$ and $F_{d}$ gates respectively.
In the multi-rail encoding, a permutation of photon paths becomes the $CN_{d}$ gate and the $d$-partite port that Fourier transforms the spatial states of photons becomes the $F_{d}$ gate and Indeed, for $d=3$ case, the multi-rail implementation of linear optical elements are given by
[…]
which can be directly generalized to any $d$-level systems. Replacing all the linear optical logic gates in Fig. 3 with the operators in (15), we obtain a circuit for the $N=3$ qutrit GHZ state in the multirail encoding as in Fig. 5.
Figure 5: $N=d=3$ GHZ generation circuit in the multirail encoding example.
An advantage of the multirail encoding circuit is that $CN_{d}$ and $\overline{CN}_{d}$ is achieved with the permutation of paths, hence the final circuit consists of simpler optical elements, i.e., $d$-partite ports including BSs, permutations of paths, and photon detectors.
Discussions.—
Our scheme has several manifest advantages in using the qudit GHZ state as a resource for high-dimensional quantum information processing. First, as we have already demonstrated, our scheme enables more efficient generations of the resource states with single-photons and linear operators, which becomes a more feasible resource state that can generate qudit cluster states with fusion gates for the qudit MBQC [23, 24, 25]. Second, our scheme is adequate to verify the optimal state for the violation of Bell-CHSH nonlocality test. While the maximally entangled state entails the maximal Bell-CHSH violation in $d=2$ systems, the scenario becomes more complicated for higher dimensional systems. Ref. [12, 13] shows that the maximal Bell-CHSH violation occurs for a non-maximally entangled state, which is also verified numerically for the $N=d=3$ case in Ref. [32]. Therefore, for the experiments of those maximal violations, schemes for generalized qudit GHZ states with relatively asymmetric amplitudes are indispensable. Our scheme easily achieve the condition by varying the transitivity of beamspliters. Third, our scheme consists of linear optical logic gates that any type of qubit encoding (e.g., optical angular momentum (OAM), multi-rail, time-bin, etc.) can realize.
In addition, the methodology of composing linear optical logic gates based on the LQG picture can be useful to find more general types of qudit multipartite entanglement. For example, Ref. [33] proposed heralded schemes for $N$-partite $N$-level symmetric and anti-symmetric states based on the method. We expect to construct more rigorous and general formalism of the LQG picture that will provide a more systematic path for the heralded generations of essential qudit entanglement such as absolutely maximally entangled (AME) states.
Acknowledgements.—
SC is grateful to Prof. Jung-Hoon Chun for his support on this research. SC is supported by National Research Foundation of Korea (NRF, 2019R1I1A1A01059964, RS-2023-00245747). JR acknowledges support from the National Research Foundation (NRF) of Korea grant funded by the Korea Government (Grant No. NRF-2022M3K2A1083890). YSK acknowledges support from the KIST institutional program (2E32941).
Appendix A Operations of $CN_{d}$, the $d$-level CNOT gate between the spatial mode and the internal state
Unlike the convential generalized CNOT gate that works on two different qudits, $CN_{d}$ and $\overline{CN}_{d}$ here transform the spatial mode of a given photon according to its internal state. They transform the two variables in the same way as the generalized CNOT gate between two qudits. The spatial mode and internal state play the role of the target and control states.
More explicitly, the operation of $CN_{d}$ given by Eqs. (Linear optical logic gates for generating $|GHZ_{N}^{d}\rangle$.—) can be enumerated as the following table:
$|\tilde{0}\rangle_{0}\to|\tilde{0}\rangle_{0}$ $|\tilde{1}\rangle_{0}\to|\tilde{1}\rangle_{1}$cdots $|\widetilde{d-1}\rangle_{0}\to|\widetilde{d-1}\rangle_{d-1}$
$|\tilde{0}\rangle_{1}\to|\tilde{1}\rangle_{1}$ $|\tilde{1}\rangle_{1}\to|\tilde{1}\rangle_{2}$cdots $|\widetilde{d-1}\rangle_{1}\to|\widetilde{d-1}\rangle_{0}$
$\vdots$ $\vdots$ $\vdots$ $\vdots$
$|\tilde{0}\rangle_{d-2}\to|\tilde{0}\rangle_{d-2}$ $|\tilde{1}\rangle_{d-2}\to|\tilde{1}\rangle_{d-1}$cdots $|\widetilde{d-1}\rangle_{d-2}\to|\widetilde{d-1}\rangle_{d-3}$
$|\tilde{0}\rangle_{d-1}\to|\tilde{0}\rangle_{d-1}$ $|\tilde{1}\rangle_{d-1}\to|\tilde{1}\rangle_{0}$cdots $|\widetilde{d-1}\rangle_{d-1}\to|\widetilde{d-1}\rangle_{d-2}$
Comparing the transformations in the same row of the table, we can see that that $CN_{d}$ splits photons of different internal states in the same spatial mode into different spatial modes.
Appendix B Comprehensive analysis on the qudit $N$-partite GHZ state generation scheme
Here we provide a comprehensive explanation on the linear optical circuit for generating the qudit GHZ state given in Fig. 2. We first explain how we designed our scheme based on the graph picture of linear quantum networks (the LQG picture) [34, 28] and then show that the scheme indeed generates the expected target state.
B.1 Construction of linear optical circuit from the boson subtraction operator
Our scheme can be considered a linear optical realization of the boson subtraction scheme in Ref. [28], which exploited the LQG picture [34, 28] to find boson subtraction schemes for generating multipartite entanglement of qubits and qudits. Those boson subtaction schemes are in principle realized as linear optical heralded circuits, which is elucidated as a set of direct translation rule for qubit cases in Ref. [35]. However, its general extension to arbitrary $d$-level cases is absent.
In Ref. [28], the bipartite graph (bigraph) that corresponds to the boson subtraction scheme for the qudit GHZ state is given by
[…]
which corresponds to the following operator (see Table I of Ref. [28]):
$\hat{A}_{N,d}=\Big{(}\frac{1}{\sqrt{2}}\Big{)}^{(d-1)N}({\hat{a}}_{1,{\tilde{0}}}-{\hat{a}}_{2,\widetilde{d-1}})^{d-1}({\hat{a}}_{2,{\tilde{0}}}-{\hat{a}}_{3,\widetilde{d-1}})^{d-1}({\hat{a}}_{3,{\tilde{0}}}-{\hat{a}}_{4,\widetilde{d-1}})^{d-1}\cdots\times({\hat{a}}_{N,{\tilde{0}}}-{\hat{a}}_{1,\widetilde{d-1}})^{d-1}.$ (16)
We can intuitively see that the following translation holds from the LQG picture to the linear optical system:
Therefore, by combing the $N$ sets of the above linear optical element with the structure of the bigraph (B.1), we obtain the scheme in Fig. 2 that generates the expected target state.
B.2 Demonstrations of the final state generated by the scheme in Fig. 2
To demonstrate that the scheme in Fig. 2 generates the qudit GHZ state, we need the following two identities that were first given in Ref. [28] without proofs. Due to their significance in calculating our circuit operation, we restate them here with proofs.
Identity 1. For $l\in\{0,1,\cdots,d-1\}$,
$({\hat{a}}_{{\tilde{0}}})^{l}({\hat{a}}_{\widetilde{d-1}})^{d-1-l}\prod_{s=0}^{d-1}{\hat{a}}^{\dagger}_{s}|vac\rangle=(-1)^{d-1-l}\frac{l!(d-1-l)!}{\sqrt{d}^{d-2}}{\hat{a}}^{\dagger}_{\widetilde{d-1-l}}|vac\rangle$ (17)
Proof.
$({\hat{a}}_{{\tilde{0}}})^{l}({\hat{a}}_{\widetilde{d-1}})^{d-1-l}\prod_{s=0}^{d-1}{\hat{a}}^{\dagger}_{s}|vac\rangle$
$=\frac{1}{\sqrt{d}^{d-1}}({\hat{a}}_{0}+{\hat{a}}_{1}+\cdots+{\hat{a}}_{d-1})^{l}({\hat{a}}_{0}+\omega{\hat{a}}_{1}+\cdots+\omega^{d-1-l}{\hat{a}}_{d-1})^{d-1-l}\prod_{s=0}^{d-1}{\hat{a}}^{\dagger}_{s}|vac\rangle$
$=\frac{1}{\sqrt{d}^{d-1}}({\hat{a}}_{0}+{\hat{a}}_{1}+\cdots+{\hat{a}}_{d-1})^{l}(d-1-l)!\sum_{\begin{subarray}{c}j_{1}\neq j_{2}\neq\cdots\\ \neq j_{d-1-l}\\ =0\end{subarray}}^{d-1}\omega^{j_{1}+j_{2}+\cdots+j_{d-1-l}}{\hat{a}}_{j_{1}}{\hat{a}}_{j_{2}}\cdots{\hat{a}}_{j_{d-1-l}}\prod_{s=0}^{d-1}{\hat{a}}^{\dagger}_{s}|vac\rangle$
$=\frac{l!}{\sqrt{d}^{d-1}}\sum_{\begin{subarray}{c}j_{1}\neq j_{2}\neq\cdots\\ \neq j_{d-1-l}\\ =0\end{subarray}}^{d-1}\omega^{j_{1}+j_{2}+\cdots+j_{d-1-l}}(\sum_{q=0}^{d-1}{\hat{a}^{\dagger}}_{q}-{\hat{a}^{\dagger}}_{j_{1}}-{\hat{a}^{\dagger}}_{j_{2}}-\cdots-{\hat{a}^{\dagger}}_{j_{d-1-l}})|vac\rangle$ (18)
Using
$\sum_{\begin{subarray}{c}j_{1}\neq j_{2}\neq\cdots\\ \neq j_{d-1-l}\\ =0\end{subarray}}^{d-1}\omega^{j_{1}+j_{2}+\cdots+j_{d-1-l}}=(-1)^{d-1-l}\sum_{j=0}^{d-1}\omega^{(d-1-l)}=0,$ (19)
Eq. (B.2) becomes
$-\frac{l!}{\sqrt{d}^{d-1}}\sum_{\begin{subarray}{c}j_{1}\neq j_{2}\neq\cdots\\ \neq j_{d-1-l}\\ =0\end{subarray}}^{d-1}\omega^{j_{1}+j_{2}+\cdots+j_{d-1-l}}({\hat{a}^{\dagger}}_{j_{1}}+{\hat{a}^{\dagger}}_{j_{2}}+\cdots+{\hat{a}^{\dagger}}_{j_{d-1-l}})|vac\rangle$
$=-\frac{l!}{\sqrt{d}^{d-1}}(d-1-l)\sum_{\begin{subarray}{c}j_{1}\neq j_{2}\neq\cdots\\ \neq j_{d-1-l}\\ =0\end{subarray}}^{d-1}\omega^{j_{1}+j_{2}+\cdots+j_{d-1-l}}{\hat{a}^{\dagger}}_{j_{1}}$
$=(-1)^{d-1-l}\frac{l!(d-1-l)!}{\sqrt{d}^{d-1}}\sum_{j_{1}=0}^{d-1}\omega^{(d-1-l)j_{1}}{\hat{a}^{\dagger}}_{j_{1}}$
$=(-1)^{d-1-l}\frac{l!(d-1-l)!}{\sqrt{d}^{d-2}}{\hat{a}}^{\dagger}_{\widetilde{d-1-l}}|vac\rangle.$ (20)
Identity 2. For $m\in\{1,\cdots,d-1\}$,
$({\hat{a}}_{{\tilde{0}}})^{m}({\hat{a}}_{\widetilde{d-1}})^{d-m}\prod_{s=0}^{d-1}{\hat{a}}^{\dagger}_{s}|vac\rangle=0$ (21)
Proof.
It is directly derived from Identity 1. ∎
There are $N$ spatial subsystems that has $d$ modes respectively in Fig. 2, hence the photon creation operator is denoted as ${\hat{a}}^{j\dagger}_{k,s}$, which denotes that a photon is in the $j$th subsystem, $k$th mode, and $s$th internal state. The $d$ photons with orthogonal qudit states are injected into each spatial subsystem. Hence the initial state is given by
$\prod_{j=1}^{N}({\hat{a}}^{j\dagger}_{0,0}{\hat{a}}^{j\dagger}_{0,1}\cdots{\hat{a}}^{j\dagger}_{0,d-1})|vac\rangle.$ (22)
Identities 1 and 2 show that when the photons pass through $CN_{d}$ gates, the only case when one photon goes to $X^{j}_{d-1-l}$ ($l\in\{1,2,\cdots,d-2\}$) is when $l$ photons goes to ${\hat{a}}^{\dagger j}_{0,\tilde{0}}$ and $(d-1-l)$ photons goes to ${\hat{a}}^{\dagger j}_{d-1,\widetilde{d-1}}$. This property can be used to verify all the surviving terms after the postselection by $D_{F}$s.
First, suppose that the $d$ photons in subsystem 1 goes to mode 0 ((${\hat{a}}^{1\dagger}_{0,\tilde{0}})^{d}$). Then by the postselections at $D_{F}$s, the only case that survives is that each set of $d$ photons in all subsystems goes to mode 0, which corresponds to the term $|\underbrace{0,0,\cdots,0}_{N}\rangle$. Second, suppose that the $d$ photons in subsystem 1 goes to mode (d-1) ((${\hat{a}}^{1\dagger}_{0,\widetilde{d-1}})^{d}$). By the same reasoning, the only surviving term under the postselection is $|\underbrace{d-1,d-1,\cdots,d-1}_{N}\rangle$. Third, suppose that among $d$ photons in subsystem 1, the $l$ photons goes to ${\hat{a}}^{1\dagger}_{0,\tilde{0}}$, $(d-1-l)$ photons to ${\hat{a}}^{1\dagger}_{d-1,\widetilde{d-1}}$, hence one photon to $X^{1}_{d-1-l}$ ($({\hat{a}}^{1\dagger}_{0,\tilde{0}})^{l}({\hat{a}}^{1\dagger}_{0,\widetilde{d-1}})^{d-1-l}{\hat{a}}^{1\dagger}_{0,\widetilde{d-1-l}}$). We can check that the surviving term to satisfy the postselection condition is $|\underbrace{d-1-l,d-1-l,\cdots,d-1-l}_{N}\rangle$. Fourth, if more than two photons go to $X^{1}_{d-1-l}$, there is no term that fulfill the postselection condition. All in all, the final state after the postselections of $D_{F}$s is given by
$\frac{1}{d^{Nd}}\Big{(}(d!)^{N}|0,0,\cdots\rangle+(d!)^{N}|d-1,d-1,\cdots,d-1\rangle+(d!\sqrt{d}\alpha)^{N}\sum_{k=1}^{d-2}|k,k,\cdots,k\rangle\Big{)}$ (23)
By setting $\alpha=\frac{1}{\sqrt{d}}$, i.e., placing $1:(d-1)$ BSs on $X^{j}_{l}$ ($j\in\{1,\cdots,N\}$, $l\in\{1,\cdots,d-2\}$), the final state becomes $|GHZ\rangle_{N}^{d}$ as expected. The success probability is given by
$P_{suc}(N,d)=d\Big{(}\frac{d!}{d^{d}}\Big{)}^{2N}.$ (24)